# Leveraging the Bias-Variance Tradeoff in Quantum Chemistry for Accurate Negative Singlet-Triplet Gap Predictions: A Case for Double-Hybrid DFT

### v1

@@ 4.2 Bias-variance tradeoff across methods
Since the typical magnitudes of negative STGs reported in various studies are $\approx 0.1$ eV or less, methods with mean errors exceeding 0.05 eV (50 meV) are generally unacceptable for drawing qualitative conclusions. Figure 2 reveals interesting trends: while some methods exhibit a high MAD, they maintain a very low SDE. These cases are highlighted in Figure 3, which plots MAD versus SDE for various methods (excluding those with SDE $>40$ meV or MAD $>250$ meV). Overall, CC3 and L-CC2 demonstrate superior performance with low MAD and SDE, whereas several DFT approaches—notably, B2GP-PLYP and PBE-QIDH—exhibit remarkably small variance. In particular, PBE-QIDH shows an SDE of less than 10 meV, indicating that its predictions are narrowly clustered around the TBE values but shifted by a constant systematic error. Although the MAD of PBE-QIDH exceeds 175 meV, systematic errors are generally easier to correct through linear regression than the non-systematic errors associated with high variance. For properties such as atomization or formation energies, DFT predictions are often corrected for systematic errors via dressed atom corrections, leading to lower errors for DFAs on the higher rungs of Jacob’s ladder [70, 56, 71].
@@
Among the local correlated methods, L-CC2 exhibits MAD and SDE values close to those of CC3, while L-ADC2 shows a slightly larger SDE than L-CC2 (Figure 3). In contrast, the dh-DFT methods B2GP-PLYP and PBE-QIDH display SDEs that are smaller than L-CC2’s, but their MADs are larger by about an order of magnitude. To explore the potential of low-variance, high-bias methods, we applied a least-squares correction to the STG predictions of B2GP-PLYP, PBE-QIDH, L-ADC(2), and L-CC2 using TBE values as the reference; the results are collected in Table 2. As expected, the MSE becomes zero after linear correction. L-ADC(2) and L-CC2 show MADs of 0.023 and 0.016 eV with SDEs of 0.017 and 0.013 eV, respectively, with only slight improvement upon correction, indicating that their errors are primarily non-systematic. In contrast, dh-DFT methods B2GP-PLYP and PBE-QIDH, which initially exhibit MADs of 0.119 and 0.177 eV and SDEs of 0.011 and 0.009 eV respectively, experience a dramatic drop in MAD (to below 0.01 eV) after correction, while retaining their low SDEs. This implies that the residual errors in these dh-DFT methods are predominantly systematic. A similar analysis for the SCS and SOS variants of these methods, presented in Table 3, confirms that all fixed dh-DFT methods show reduced MAD and SDE values, indicating that their prediction errors are largely systematic when compared to TBE.
@@
Table 2: For various methods (aug-cc-pVDZ basis set) STGs of benchmark systems (Figure 1) are presented before and after bias-correction. Error metrics and the reference TBE values are also provided. All values are in eV.
# B2GP-PLYP PBE-QIDH L-ADC2 L-CC2 TBE
@@
Table 3: Error metrics for predicting twelve values of S1 & T1 energetics of triangular molecules reported in Ref. [14]. Values are reported compared to the theoretical best estimates (TBE) from Ref. [14]. In all cases, the basis set is aug-cc-pVDZ, and the geometries are from Ref. [14]. MSE: mean signed error, MAE: mean absolute error, SDE: standard deviation of the error, minE: minimal error, and maxE: maximal error. Wherever necessary, the convention for deviation is ‘Method--TBE.’ All values are in eV.
@@
B2GP-PLYP S1 1.0350 -0.2382 0.184 0.030 0.012 0.020
T1 1.0294 -0.1105 0.065 0.031 0.018 0.024
STG 1.0654 -0.1190 0.119 0.011 0.007 0.009
SOS-B2GP-PLYP21 S1 1.0164 -0.2574 0.231 0.015 0.008 0.011
T1 1.0048 -0.1889 0.181 0.019 0.013 0.018
STG 0.9167 -0.0561 0.050 0.013 0.007 0.010
SCS-B2GP-PLYP21 S1 1.0086 -0.2948 0.281 0.013 0.009 0.012
T1 1.$0$0.2164 0.216 0.020 0.015 0.020
STG 0.9246 -0.0685 0.064 0.012 0.008 0.009
PBE-QIDH S1 0.9627 -0.2158 0.277 0.034 0.013 0.023
T1 0.9693 -0.0514 0.100 0.033 0.017 0.025
STG 1.0284 -0.1787 0.177 0.009 0.008 0.009
SOS-PBE-QIDH S1 1.0120 -0.1281 0.110 0.012 0.008 0.009
T1 0.9889 -0.0862 0.104 0.017 0.012 0.016
STG 0.8423 -0.0246 0.018 0.021 0.010 0.012
SCS-PBE-QIDH S1 1.0091 -0.1479 0.134 0.011 0.007 0.009
T1 0.9907 -0.0818 0.096 0.018 0.013 0.017
STG 0.8815 -0.0477 0.038 0.017 0.009 0.011
  • a Calculated with STG calculated using linearly corrected S1 and T1 values.
  • b Calculated using linearly corrected STG values.
@@ 4.3 Internally referenced bias-variance correction for double-hybrid DFT
Here, the coefficients $a_{x}$ and $a_{c}$ control the fraction of exact exchange (via Hartree–Fock) and MP2-level correlation, respectively. In the standard formulation, PBE-QIDH employs $a_{x}=0.69$ and $a_{c}=0.33$ [72] and belongs to a family of dh-DFAs with $a_{c}=a_{x}^{3}$, which have been found to deliver low errors for STG [73]. Similarly, B2GP-PLYP is based on the GGA-B88 exchange functional and a combination of LYP-GGA and VWN3-LDA correlation functionals:
@@
Figure 4: Dependence of prediction errors of dh-DFT with exchange and correlation mixing parameters $a_{x}$ and $a_{c}$ (See Eq. 4.1 and Eq. 4.2). MSE, MAD and SDE (in meV) are shown B2GP-PLYP (left panels) and PBE-QIDH (right panels) compared to TBE. In all plots, the default mixing coefficients are marked by a red cross—(0.65, 0.36) for B2GP-PLYP and (0.69, 0.33) for PBE-QIDH—and coefficients that minimize MAD are marked by a blue star—(0.75, 0.55) for both B2GP-PLYP and PBE-QIDH.
@@
Using B2GP-PLYP and PBE-QIDH DFAs, we scanned the $a_{x}$ and $a_{c}$ values and computed error metrics for all 12 benchmark STGs using TBE as the reference. Some calculations failed to converge when $a_{x}$ and $a_{c}\approx 0$. As shown in Figure 4, both dh-DFAs exhibit low SDEs for the default mixing parameters, and a continuous range of ($a_{x}$, $a_{c}$) combinations yielding zero MSE is observed (Figure 4, top panels). However, the minimum MAD is achieved at $a_{x}=0.75$ and $a_{c}=0.55$, with both DFAs delivering MADs of approximately 25 meV compared to TBE. We denote these optimized parameterizations as B2GP-PLYP (75,55) and PBE-QIDH (75,55). To assess the variation of optimal $a_{x}$ and $a_{c}$ across error metrics for S1 and T1 energies separately, we performed a comparative analysis (Table S1) using PBE-QIDH. Notably, increased accuracy in S1 energies is observed when both mixing parameters approach 1 (with $a_{x}=0.90$ and $a_{c}=0.85$ yielding a MAD of 68 meV), while the lowest MAD for T1 (22 meV) is obtained for $a_{x}=0.5$ and $a_{c}=0.15$.
@@
Figure 5: Comparison of 12 STGs of benchmark systems across methods: a) Scatterplot of values predicted by various methods (with the aug-cc-pVDZ basis set) versus theoretical best estimates (TBE). b) Distribution of errors in predicted STGs relative to TBE shown as smooth densities.
@@ 4.4 Application to heptazine derivatives
To probe the applicability of our bias-correction strategy, we focused on four DFIST candidates proposed in Ref. [4]. At the T1 geometry, the STGs of HzTFEX2, HzPipX2, HzTFEP2, and HzTFET2 at the EOM-CCSD, ADC(2), L-CC2, and SCS-ADC(2) levels exhibit smaller magnitudes[4] than those of the twelve triangular systems discussed above. This difference is attributed to the change in geometry, as STGs generally increase when moving away from the ground state minimum (S0) as pointed out from the context of excited-state nuclear dynamics[77] and the pseudo-Jahn–Teller effect [25]. Despite the larger molecular size of the heptazine derivatives compared to the triangular systems, local methods such as L-ADC(2) and L-CC2 provide significant speedups over their canonical counterparts, while dh-DFT calculations remain computationally feasible. This enables a straightforward application of our internally referenced scaling procedure, which corrects systematic errors without requiring external high-level reference data.
@@
For all four systems, geometry optimization was performed at the $\omega$B97X-D3/def2TZVP level (see SI for details), followed by single-point excited-state energy calculations. The STGs of these four systems, obtained from various methods, are presented in Figure 6. In the absence of a high-fidelity reference such as theoretical best estimates for these systems, L-CC2 results are considered the most reliable. We used B2GP-PLYP (75,55) energies as references to scale the predictions from B2GP-PLYP (65,36) and similarly corrected PBE-QIDH (69,33) predictions using the (75,55) configuration. Notably, the application of the internal scaling does not require any additional high-level reference to identify the low-bias or low-variance domain in the parameter space of dh-DFAs.
Among the results from various dh-DFAs, SCS-PBE-QIDH shows the best agreement with L-CC2, achieving a MAD of 0.014 eV. In contrast, bias-corrected results for B2GP-PLYP and PBE-QIDH are systematically more negative than the L-CC2 values, with MADs of 0.057 eV and 0.048 eV, respectively. While one might argue that simply shifting the results of these two DFAs could render the STGs negative, our correction is performed without requiring reference values from another method. As in the case of triangulene systems, B2GP-PLYP exhibits a smaller bias compared to PBE-QIDH, as reflected in its predicted STGs of --0.132, --0.042, --0.107, and --0.105 eV. Furthermore, we observed that L-CC2’s error relative to the TBE is consistently shifted by -0.013 eV, whereas the errors of the (75,55) variant of the two dh-DFAs are centered around zero (albeit with a large spread), as shown in Figure 5. Consequently, compared to L-CC2, the bias-corrected dh-DFAs are expected to exhibit a non-vanishing MSE; indeed, the MSE values for B2GP-PLYP (75,55) and PBE-QIDH (75,55) are -0.057 eV and -0.048 eV, respectively. Compared to a more accurate reference such as CC3 (which is practically challenging for the four heptazine derivatives), our bias-corrected results may show better agreement than the other methods discussed. Moreover, it is possible to identify combinations of $a_{x}$ and $a_{c}$ that yield low bias and low variance relative to L-CC2, which can be further explored for extending the bias correction scheme to larger datasets lacking high-level reference data.
@@ 5 Conclusions
First principles excited state modeling plays a crucial rule in the identification of new molecular light emitters with a negative S1-T1 energy gap. Historically, theoretical interest in such molecules existed for a long time in the context of Hund’s rule violation[78, 79, 80, 81, 82]. Yet, revived interest in them increased since their identification through first principles modeling of heptazine[83, 84]. Although individual S1 and T1 energies are challenging to predict with high precision, the error cancellation inherent in STG calculations allows for improved accuracy, even when the reported gaps are less than 0.1 eV. Benchmarking twelve STG energy gaps in triangular molecules, we find that local wavefunction methods—specifically L-CC2 and L-ADC2—deliver mean errors of approximately 15 meV with correspondingly low standard deviations, providing both exceptional accuracy and the computational speed required for high-throughput screening of medium size molecules such as azaphenalenes with bulky substitutions.
In parallel, we show that double-hybrid DFT approximations—though initially marked by high systematic errors—can be significantly improved through linear regression-based bias correction without depending on a high-level reference. For example, the PBE-QIDH method in its default configuration (69% exact exchange and 33% MP2-level correlation) exhibits an impressively low variance (9 meV) but high mean errors; by adjusting the parameters to 75% exchange and 55% correlation, the mean error is reduced to 4 meV, albeit with increased variance. When applying this strategy to four heptazine derivatives, the SCS and SOS variants of both B2GP-PLYP and PBE-QIDH—though not subjected to bias correction—yield STG predictions in better agreement with L-CC2, suggesting that these variants naturally mitigate some of the bias in standard dh-DFT. However, bias correction remains valuable for further refining low-variance models, particularly when an appropriate low-bias reference is available.
The purpose of this study is not to advocate different models for different properties but to leverage the intrinsic bias-variance tradeoff for accurate STG predictions. By using a suitable low-bias method as an internal reference, our approach corrects systematic errors in low-variance models, thereby enabling quantitatively reliable negative STG estimations. Open questions remain regarding the performance of these parameterized models for positive STGs and systems with charge-transfer excitations[85], where even methods like CC2 may exhibit deficiencies. Robust application of our strategy requires reliable reference datasets and could benefit from further bias-variance analysis using alternative information-based metrics [86]. Overall, our findings establish a mathematically sound framework for predictive modeling in first-principles molecular design, with promising implications for the development of novel light emitters and other applications.

### v2

@@ 4.2 Bias-variance tradeoff across methods
If the magnitude of negative STGs exceeds 0.2 eV, methods with mean errors $\approx 0.1$ eV may still allow for qualitative identification. However, for systems where the STG is below $\approx 0.1$ eV, similar prediction errors become too large to yield reliable qualitative conclusions. Figure 2 reveals interesting trends in the bias-variance tradeoff in various methods: while some methods exhibit a high MAD, they maintain a very low SDE. These cases are highlighted in Figure 3, which plots MAD versus SDE for various methods (excluding those with high MAD or SDE). Overall, CC3 and L-CC2 demonstrate superior performance with low MAD and SDE, whereas several DFT approaches—notably, B2GP-PLYP and PBE-QIDH—exhibit remarkably small variance. In particular, PBE-QIDH shows an SDE of less than 10 meV, indicating that its predictions are narrowly clustered around the TBE values but shifted by a constant systematic error. Although the MAD of PBE-QIDH exceeds 175 meV, systematic errors are generally easier to correct through linear regression than the non-systematic errors associated with high variance. For properties such as atomization or formation energies, DFT predictions are often corrected for systematic errors via dressed atom corrections, leading to lower errors for DFAs on the higher rungs of “Jacob’s ladder” [70, 56, 71].
@@
Among the local correlated methods, L-CC2 exhibits MAD and SDE values close to those of CC3, while L-ADC2 shows a slightly larger SDE than L-CC2 (Figure 3). In contrast, the dh-DFT methods B2GP-PLYP and PBE-QIDH display SDEs that are smaller than L-CC2’s, but their MADs are larger by about an order of magnitude. To explore the potential of low-variance, high-bias methods, we applied a least-squares correction to the STG predictions of B2GP-PLYP, PBE-QIDH, L-ADC(2), and L-CC2 using TBE values as the reference; the results are collected in Table 2. As expected, the MSE becomes zero after linear correction. L-ADC(2) and L-CC2 show MADs of 0.023 and 0.016 eV with SDEs of 0.017 and 0.013 eV, respectively, with only slight improvement upon correction, indicating that their errors are primarily non-systematic. In contrast, dh-DFT methods B2GP-PLYP and PBE-QIDH, which initially exhibit MADs of 0.119 and 0.177 eV and SDEs of 0.011 and 0.009 eV respectively, experience a dramatic drop in MAD (to below 0.01 eV) after correction, while retaining their low SDEs. This implies that the errors in these dh-DFT methods are predominantly systematic. A similar analysis for the SCS and SOS variants of these methods, presented in Table 3, confirms that all fixed dh-DFT methods show reduced MAD and SDE values, indicating that their prediction errors are largely systematic when compared to TBE.
@@
Table 2: For various methods (aug-cc-pVDZ basis set) STGs of benchmark systems (Figure 1) are presented before and after bias-correction. Error metrics and the reference TBE values are also provided. MSE: mean signed error, MAE: mean absolute error, and SDE: standard deviation of the error. Wherever necessary, the convention for deviation is ‘Method--TBE.’ All values are in eV.
# B2GP-PLYP PBE-QIDH L-ADC2 L-CC2 TBEe
@@
Table 3: Error metrics for predicting twelve values of S1 and T1 energetics of triangular molecules reported in Ref. [14]. Values are reported compared to the theoretical best estimates (TBE) from Ref. [14]. In all cases, the basis set is aug-cc-pVDZ, and the geometries are from Ref. [14]. MSE: mean signed error, MAE: mean absolute error, and SDE: standard deviation of the error. Wherever necessary, the convention for deviation is ‘Method--TBE.’ All values are in eV.
@@
B2GP-PLYP S1 1.0350 --0.2382 0.184 0.030 0.012 0.020
T1 1.0294 --0.1105 0.065 0.031 0.018 0.024
STG 1.0654 --0.1190 0.119 0.011 0.007 0.009
SOS-B2GP-PLYP21 S1 1.0164 --0.2574 0.231 0.015 0.008 0.011
T1 1.0048 --0.1889 0.181 0.019 0.013 0.018
STG 0.9167 --0.0561 0.050 0.013 0.007 0.010
SCS-B2GP-PLYP21 S1 1.0086 --0.2948 0.281 0.013 0.009 0.012
T1 1.$0$-0.2164 0.216 0.020 0.015 0.020
STG 0.9246 --0.0685 0.064 0.012 0.008 0.009
PBE-QIDH S1 0.9627 --0.2158 0.277 0.034 0.013 0.023
T1 0.9693 --0.0514 0.100 0.033 0.017 0.025
STG 1.0284 --0.1787 0.177 0.009 0.008 0.009
SOS-PBE-QIDH S1 1.0120 --0.1281 0.110 0.012 0.008 0.009
T1 0.9889 --0.0862 0.104 0.017 0.012 0.016
STG 0.8423 --0.0246 0.018 0.021 0.010 0.012
SCS-PBE-QIDH S1 1.0091 --0.1479 0.134 0.011 0.007 0.009
T1 0.9907 --0.0818 0.096 0.018 0.013 0.017
STG 0.8815 --0.0477 0.038 0.017 0.009 0.011
@@ 4.3 Internally referenced bias-variance correction for double-hybrid DFT
Here, the coefficients $a_{x}$ and $a_{c}$ control the fraction of exact exchange (via Hartree–Fock) and MP2-level correlation, respectively. In the standard formulation, PBE-QIDH ($a_{x}=0.69$ and $a_{c}=0.33$)[72] belongs to a family of dh-DFAs with $a_{c}=a_{x}^{3}$ delivering low errors for STG [73]. Similarly, B2GP-PLYP is based on the GGA-B88 exchange functional and a combination of LYP-GGA and VWN3-LDA correlation functionals:
@@
Figure 4: Dependence of prediction errors of dh-DFT with exchange and correlation mixing parameters $a_{x}$ and $a_{c}$ (See Eq. 4.1 and Eq. 4.2). MSE, MAD and SDE (in meV) are shown for B2GP-PLYP (left panels) and PBE-QIDH (right panels) compared to TBE. Values of $a_{x}$ and $a_{c}$ were sampled in steps of 0.05 and the surfaces were smoothened using cubic functions. In all plots, the default mixing coefficients are marked by a red cross—(0.65, 0.36) for B2GP-PLYP and (0.69, 0.33) for PBE-QIDH—and coefficients that minimize MAD are marked by a blue star—(0.75, 0.55) for both B2GP-PLYP and PBE-QIDH.
@@
Using B2GP-PLYP and PBE-QIDH DFAs, we scanned $a_{x}$ and $a_{c}$, and computed error metrics for all 12 benchmark STGs using TBE as the reference. As shown in Figure 4, both dh-DFAs exhibit low SDEs for the default mixing parameters, and a continuous range of ($a_{x}$, $a_{c}$) combinations yielding zero MSE is observed (Figure 4, top panels). Some calculations failed to converge when $a_{x}$ and $a_{c}\approx 0$; these are not shown in Figure 4. However, the minimum MAD is achieved at $a_{x}=0.75$ and $a_{c}=0.55$, with both DFAs delivering MADs of approximately 25 meV compared to TBE. We denote these optimized parameterizations as B2GP-PLYP (75,55) and PBE-QIDH (75,55). To assess the variation of optimal $a_{x}$ and $a_{c}$ across error metrics for S1 and T1 energies separately, we performed a comparative analysis (Table S1) using PBE-QIDH. Notably, increased accuracy in S1 energies is observed when both mixing parameters approach 1 (with $a_{x}=0.90$ and $a_{c}=0.85$ yielding a MAD of 68 meV), while the lowest MAD for T1 (22 meV) is obtained for $a_{x}=0.5$ and $a_{c}=0.15$.
@@
Figure 5: Comparison of 12 STGs of benchmark systems across methods: a) Scatterplot of values predicted by various methods (with the aug-cc-pVDZ basis set) versus theoretical best estimates (TBE). TBE and CC3 results are from Ref. [14]. b) Distribution of errors in predicted STGs relative to TBE shown as smooth densities.
@@ 4.4 Application to heptazine derivatives
To probe the applicability of our bias-correction strategy, we focused on four DFIST candidates proposed in Ref. [4]. At the T1 geometry, the STGs of HzTFEX2, HzPipX2, HzTFEP2, and HzTFET2 at the EOM-CCSD, ADC(2), L-CC2, and SCS-ADC(2) levels exhibit smaller magnitudes[4] than those of the twelve triangular systems discussed above. This difference is attributed to the change in geometry, as STGs generally increase when moving away from the ground state minimum (S0) as pointed out from the context of excited-state nuclear dynamics[77] and the pseudo-Jahn–Teller effect [25]. Despite the larger molecular size of the heptazine derivatives compared to the triangular systems, local methods such as L-ADC(2) and L-CC2 provide significant speedups over their canonical counterparts, while dh-DFT calculations remain computationally feasible.
@@
For all four systems, geometry optimization was performed at the $\omega$B97X-D3/def2TZVP level (see SI for details), followed by single-point excited-state energy calculations. The STGs of these four systems, obtained from various methods, are presented in Figure 6. In the absence of a high-fidelity reference such as TBE for these systems, L-CC2 results are considered the most reliable. We used B2GP-PLYP (75,55) energies as references to scale the predictions from B2GP-PLYP (65,36) and similarly corrected PBE-QIDH (69,33) predictions using the (75,55) configuration. Notably, the application of the internal scaling does not require any additional high-level reference to identify the low-bias or low-variance domain in the parameter space of dh-DFAs. This enables a straightforward application of our internally referenced scaling procedure, which corrects systematic errors without requiring external high-level reference data.
Among the results from various dh-DFAs, SCS-PBE-QIDH shows the best agreement with L-CC2, achieving an MAD of 0.014 eV. In contrast, bias-corrected results for B2GP-PLYP and PBE-QIDH are systematically more negative than the L-CC2 values, with MADs of 0.057 eV and 0.048 eV, respectively. While one might argue that simply shifting the results of these two DFAs could render the STGs negative, our correction is performed without requiring reference values from another method. As in the case of triangular benchmark systems (shown in Figure1), B2GP-PLYP exhibits a smaller bias compared to PBE-QIDH. Furthermore, for triangular systems, we observed that L-CC2’s error relative to the TBE is consistently shifted by --0.013 eV, whereas the errors of the (75,55) variant of the two dh-DFAs are centered around zero (albeit with a large spread), as shown in Figure 5. Consequently, compared to L-CC2, the bias-corrected dh-DFAs are expected to exhibit a non-vanishing MSE; indeed, the MSE values for B2GP-PLYP (75,55) and PBE-QIDH (75,55) are --0.057 eV and --0.048 eV, respectively. Compared to a more accurate reference such as CC3 (which is practically challenging for the four heptazine derivatives), our bias-corrected results may show better agreement than the other methods discussed. Moreover, it is possible to identify combinations of $a_{x}$ and $a_{c}$ that yield low bias and low variance relative to L-CC2, which can be further explored for extending the bias correction scheme to larger datasets lacking high-level reference data.
@@ 5 Conclusions
First principles excited state modeling plays a crucial rule in the identification of new molecular light emitters with a negative S1-T1 energy gap. Historically, theoretical interest in such molecules existed for a long time in the context of Hund’s rule violation[78, 79, 80, 81, 82]. Yet, revived interest in them increased since their identification through first principles modeling[83, 84]. Although individual S1 and T1 energies are challenging to predict with high precision, the error cancellation inherent in STG calculations allows for improved accuracy, even when the reported gaps are less than 0.1 eV. Benchmarking twelve STG energy gaps in triangular molecules, we find that local wavefunction methods—specifically L-CC2 and L-ADC2—deliver mean errors of approximately 15 meV with correspondingly low standard deviations, providing both exceptional accuracy and the computational speed required for high-throughput screening of medium size molecules such as azaphenalenes with bulky substitutions.
In parallel, we show that double-hybrid DFT approximations—though initially marked by high systematic errors—can be significantly improved through linear regression-based bias correction without depending on a high-level reference. For example, the PBE-QIDH method in its default configuration (69% exact exchange and 33% MP2-level correlation) exhibits an impressively low variance (9 meV) but high mean errors; by adjusting the parameters to 75% exchange and 55% correlation, the mean error is reduced to 4 meV, albeit with increased variance. We propose a bias correction strategy to leverage the bias-variance tradeoff intrinsic the dh-DFA formalisms. Using a suitable low-bias method as an internal reference, the approach corrects systematic errors in low-variance models, thereby enabling quantitatively reliable negative STG estimations. In the case of four heptazine derivatives, the SCS and SOS variants of both B2GP-PLYP and PBE-QIDH—though not subjected to bias correction—yield STG predictions in better agreement with L-CC2, suggesting that these variants naturally mitigate some of the bias in standard dh-DFT. However, bias correction remains valuable for further refining low-variance models, particularly when an appropriate low-bias reference is available.
The purpose of this study is not to advocate different models for different properties but to leverage the intrinsic bias-variance tradeoff for accurate STG predictions. Open questions remain regarding the performance of these parameterized models for positive STGs and systems with charge-transfer excitations[85]. Robust application of our strategy requires reliable reference datasets and could benefit from further bias-variance analysis using alternative information-based metrics [86]. Overall, the findings of this study establish a mathematically sound framework for predictive modeling in first-principles molecular design, with promising implications for the development of novel light emitters and other applications.